OpenStudy (anonymous):

MEDAL WILL BE REWARDED What is the product in simplest form? State any restrictions on the variable.

OpenStudy (anonymous):



OpenStudy (anonymous):

is that circle thing supposed to be a multiplication thing?

OpenStudy (anonymous):

yea sorry lol kinda looks funny

OpenStudy (anonymous):

\[\frac{ y^2 }{ y-3 }\times \frac{ y^2-y-6 }{ y^2+y }\] there that looks better lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So factor the right-hand rational factor into \[\frac{ (y-3)(y+2) }{ y(y+1) }\] Now cancel common factors on top and bottom. This gets rid of a factory of y, and a factor of (y-3), leaving only \[\frac{y(y+2)}{y+1}\] The original expression would not be defined if y were 3, 0, or -1, so this is defined for all values except these three.

OpenStudy (anonymous):

@AddemF not going to lie you lost me

OpenStudy (anonymous):

what did you get lost by?

OpenStudy (anonymous):

Well, do you understand how I factored the right-hand rational expression?

OpenStudy (jennychan12):

i factored out the common stuff \[\frac{ y^2 }{ y-3 }*\frac{ (y-3)(y+2) }{ y(y+1) }\]  final answer is \[\frac{ y(y+2) }{ y+1 } = \frac{ y^2+2y }{ y+1 }\]